OpenStudy (anonymous):

cos^2 (1/2sin^-1 3/5)

myininaya (myininaya):

try this \[\text{ Let } sin(u)=\frac{3}{5}\] use to building a right triangle You can now find cos and then you should be able to find \[\cos^2(\frac{1}{2}u)\] using half angle identity and the triangle